# FINITE-DIMENSIONAL GAUSSIAN APPROXIMATION FOR DEEP NEURAL NETWORKS: UNIVERSALITY IN RANDOM WEIGHTS

### v1

@@ Abstract
We study the Finite-Dimensional Distributions (FDDs) of deep neural networks with randomly initialized weights that have finite-order moments. Specifically, we establish Gaussian approximation bounds in the Wasserstein-$1$ norm between the FDDs and their Gaussian limit assuming a Lipschitz activation function and allowing the layer widths to grow to infinity at arbitrary relative rates. In the special case where all widths are proportional to a common scale parameter $n$ and there are $L-1$ hidden layers, we obtain convergence rates of order $n^{-({1}/{6})^{L-1}+\epsilon}$, for any $\epsilon>0$.
@@ 1 Introduction
Apollonio et al. (2025) non-Lip FDD $W_{1}$ (w.r.t. 2-norm) & Convex distance $n^{-\tfrac{1}{2}}$
Trevisan (2023) Lip FDD $W_{p}$ (w.r.t. $p$-norm) for $p\in[1,\infty)$ $n^{-\tfrac{1}{2}}$ $n^{-1}$ (full-rank limiting covariance)
@@
Celli and Peccati (2025) Piecewise linear FDD TV & $W_{2}$ (w.r.t. 2-norm) $n^{-1}$ (full-rank limiting covariance)
@@
where $\sigma$ is applied coordinate-wise. Typically, each layer also includes bias parameters; however, setting them to zero does not affect our results and is therefore omitted—see Remark 1.5. Our main result establishes Gaussian approximation bounds, in the Wasserstein-1 distance, between the finite-dimensional distributions (FDDs) of wide neural networks and their Gaussian process limits, under general weight distributions satisfying mild moment conditions and assuming a Lipschitz activation function.
@@
Neal (1996) showed that the output distribution of a single hidden-layer neural network converges to a Gaussian in the infinite-width limit. This is particularly straightforward to notice because $F^{(2)}$ can be written as a sum of independent functions and so classical techniques for sums of independent random variables immediately result in Gaussian behavior. Subsequent works such as Lee et al. (2018) and de G. Matthews et al. (2018) provided heuristic arguments and empirical evidence suggesting that deep neural networks with multiple layers similarly exhibit Gaussian behavior as their widths grow. More recently, Hanin (2023) rigorously established asymptotic convergence results for deep networks, further supporting the Gaussian behavior in the infinite-width regime.
@@
Apart from their intrinsic probabilistic interest, there are several compelling reasons to study deep neural networks (DNNs) with non-Gaussian weights. In practice, weights are typically initialized randomly—often with independent entries—and subsequently optimized using algorithms like stochastic gradient descent (SGD) to minimize a loss function over a given dataset. While Gaussian initializations are widely used, other schemes such as uniform initialization (Glorot and Bengio, 2010; Golikov and Yang, 2022) and Bernoulli distributions, particularly in quantized networks, (Li et al., 2017) are also common. Moreover, in transfer learning settings where pre-trained models are fine-tuned on new tasks, the initial weight distribution is often far from Gaussian. This motivates the study of DNN behavior under more general random weight initializations. Furthermore, Jacot et al. (2018) showed that in the so-called lazy training regime, infinitely wide DNNs exhibit deterministic training dynamics governed by the Neural Tangent Kernel (NTK), and their fluctuations converge to Gaussian processes—emphasizing how the choice of initialization fundamentally shapes both the limiting kernel and the resulting generalization behavior (Carvalho et al., 2023). Finally, several works have explored DNNs with weights drawn from heavy-tailed or stable distributions, which may exhibit infinite variance (Der and Lee, 2005; Favaro et al., 2023; Bordino et al., 2023; Soto, 2024; Jung et al., 2023; Loría and Bhadra, 2023). These studies are motivated by empirical advantages observed in such models and have established that, in the infinite-width limit, network outputs may converge to stable processes rather than Gaussian processes.
@@
A recent and very thorough account of this literature is given in (Celli and Peccati, 2025, Section 1.7), but we give a quick review to contextualize our results. A summary of quantitative bounds for deep neural networks is provided in Table 1. For shallow NNs, some functional results include Eldan et al. (2021), Klukowski (2022), and Cammarota et al. (2024), all with fairly restrictive assumptions on the activations or weight distributions, and typically in weak $L^{p}$-type metrics at the functional level. For deep NNs, rates of convergence for FDDs assuming Gaussian weights have been given in Apollonio et al. (2025), Basteri and Trevisan (2024), Bordino et al. (2024), Celli and Peccati (2025), Favaro et al. (2025), and Trevisan (2023), under a variety of activations and metrics. At the functional level, assuming Gaussian weights, Favaro et al. (2025) derive rates of convergence with respect to Sobolev topologies in the Wasserstein-$2$ norm, and, assuming the activation is infinitely smooth, with respect to the supremum topology. Assuming general weights, process convergence with respect to the supremum topology was shown by Hanin (2023), with rates of convergence in the same topology in the Wasserstein-$1$ norm given by Balasubramanian et al. (2024).
@@ 1.1 Main result
In this work, we provide a bound on the Wasserstein-$1$ norm between the FDDs of $F^{(L)}$ and their Gaussian limit assuming a general weight distribution and a Lipschitz activation function. The bound is completely explicit (albeit complicated) in all parameters and tends to zero as the widths tend to infinity in any way, for fixed $L$. This is the first such bound appearing in the literature; see also Table 1.
@@
Given these, the limiting Gaussian field is defined recursively via intermediate fields $G^{(\ell)}$, $\ell=1,\ldots,L-1$ as follows. We first set $G^{(1)}=F^{(1)}$, and, given we have defined the distribution of $G^{(\ell)}$, we let $G^{(\ell+1)}$ be a Gaussian random field with covariance
$C_{ij}^{(\ell+1)}(x,y)=\delta_{ij}\biggl{(}c_{w}^{(\ell)}\mathop{{}\mathbbm{E}}\mathopen{}\Bigl{[}\sigma\bigl{(}G^{(\ell)}_{1}(x)\bigr{)}\sigma\bigl{(}G^{(\ell)}_{1}(y)\bigr{)}\Bigr{]}\biggr{)},$
where $\delta_{ij}$ denotes the Kronecker delta. Note that the covariance function of $F^{(L)}$, calculated from (1.1) and (1.2), follows the same recursion, but with $G$ replaced by $F$, so that the coordinates of $F^{(\ell)}$ are uncorrelated, though not independent. An important component of most proofs in the deep regime is that if $F^{(\ell)}$ is approximately Gaussian, then the coordinates should be “close” to independent.
@@
$\mathsf{d}_{1}\bigl{(}F^{(L)}(\chi),G^{(L)}(\chi)\bigr{)}:=\sup_{\zeta\in\mathcal{W}_{n_{L}\times s}}\bigl{\lvert}\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}\zeta(F^{(L)}(\chi))\bigr{]}-\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}\zeta(G^{(L)}(\chi))\bigr{]}\bigr{\rvert},$
@@
Retaining the notation above, assume there is $p>2$ and constants $c_{2p}^{(\ell)}\geqslant 1$ for $\ell=0,\ldots,L-2$ such that $\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}(W_{ij}^{(\ell)})^{2p}\bigr{]}\leqslant c_{2p}^{(\ell)}/n_{\ell}^{p}$ and a constant $c_{3}^{(L-1)}$ such that $\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}\lvert W_{ij}^{(L-1)}\rvert^{3}\bigr{]}\leqslant c_{3}^{(L-1)}n_{L-1}^{-3/2}$, Then there is a constant $\mathtt{C}$ depending on $\sigma$, $p$, $L$, $\chi$, $c_{2p}^{(\ell)}$, $\ell=0,\ldots,L-2$ and $c_{3}^{(L-1)}$ such that
$\mathsf{d}_{1}\bigl{(}F^{(L)}(\chi),G^{(L)}(\chi)\bigr{)}\leqslant\mathtt{C}\,n_{L}^{1/3}\sum_{m=1}^{L-1}n_{m}^{-\frac{1}{6}\left(\frac{p-2}{3(2p-1)}\right)^{L-m-1}}.$
@@
$n_{L}^{1/3}\sum_{m=1}^{L-1}\frac{1}{n_{m}^{\left(1/6\right)^{L-m}-\delta}},$ (1.3)
@@
$n_{L}^{\varphi}\sum_{m=1}^{L-1}n_{m}^{-\frac{1}{2}},$
for some $\varphi>0$. Our worse power stems from initially working in an integral probability metric $\mathsf{d}_{3}$ defined at (2.1), which is weaker than $\mathsf{d}_{1}$, but can be easily used with Stein’s method and does not require any assumptions on the covariance matrix of the limiting Gaussian vector $G^{(L)}(\chi)$ (see the forthcoming Remark 1.4), without which the rate may in truth be worse than the classical setting. We then use a smoothing argument to get back to $\mathsf{d}_{1}$ which introduces a factor of $1/3$ to the power. (There is a further factor of $1/2$ introduced for reasons too technical to explain in a remark, but it stems from (2.13).) The argument is inductive in the layers, and the smoothing is performed at each step, which leads to the power of $(L-m)$ in the factor of $1/6$ in the power. Similar to (Balasubramanian et al., 2024, Section 6.2), the rate could be somewhat improved by assuming the activation function has three bounded derivatives. The form of the rate would still have the power of $(L-m)$ in the exponent, but now with a base of $(1/2)$ (stemming from (2.13)) rather than $(1/6)$.
@@
It is worth noting that our result requires no assumption on the limiting covariance of $G^{(L-1)}(\chi)$. As previously mentioned, this is because we first work in the weaker $\mathsf{d}_{3}$ metric; see Theorem 2.10. The potential degeneracy of the covariance figures prominently in Gaussian approximation results for DNNs, and many rates depend on the eigenvalues of the limiting covariance, including demanding they are all strictly positive (see Table 1). Applying such results require additional hypotheses on the NN, and $\sigma$ especially. In contrast, our constants are totally explicit and do not involve eigenvalues of the covariance. Revisiting the discussion of the best achievable rates of Remark 1.3, since DNNs do not fit neatly into the setting of the classical central limit theorem, the rate (1.3) may not be achievable without additional assumptions.
@@
As is typical with approximation results for DNNs in transportation metrics, the biases that usually appear in the definition (1.1) of the NN do not affect the error. See (Balasubramanian et al., 2024, Proof of Theorem 1.2) for the argument in the more challenging infinite-dimensional setting.
@@ 2 The proof
$\mathsf{d}_{3}\bigl{(}F^{(L)}(\chi),G^{(L)}(\chi)\bigr{)}:=\sup_{\zeta\in\mathcal{F}_{n_{L}\times s}}\bigl{\lvert}\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}\zeta(F^{(L)}(\chi))\bigr{]}-\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}\zeta(G^{(L)}(\chi))\bigr{]}\bigr{\rvert},$ (2.1)
@@
where $\partial_{(\cdot)}$ denotes the partial derivative with respect to the coordinate(s) given by ()($\cdot$)( ⋅ ). We then move back to the Wasserstein distance using standard smoothing arguments; see Lemma 2.11.
For $F^{(\ell)}\in\mathbbm{R}^{n_{\ell}\times s}$ and $1\leqslant d\leqslant n_{\ell}$, write $F_{[d]}^{(\ell)}:=(F^{(\ell)}_{1},\ldots,F^{(\ell)}_{d})\in\mathbbm{R}^{d\times s}$. Our first step is to look at the distance between an intermediate field. Let $(\widetilde{W}^{(\ell)})_{\ell=0}^{L-1}$ be Gaussian i.i.d. weight matrices having the same variances as $(W^{(\ell)})_{\ell=0}^{L-1}$ given at (1.2), with the two sequences of weight matrices independent so that, e.g., $\widetilde{W}^{(\ell-1)}$ is independent of $F^{(\ell-1)}$. Let $\widetilde{F}^{(\ell)}=\widetilde{W}^{(\ell-1)}\sigma(F^{(\ell-1)})$. Then the triangle inequality implies for an $d=1,\ldots,n_{L}$, we have

### v2

@@ Abstract
We study the Finite-Dimensional Distributions (FDDs) of deep neural networks with randomly initialized weights that have finite-order moments. Specifically, we establish Gaussian approximation bounds in the Wasserstein-$1$ norm between the FDDs and their Gaussian limit assuming a Lipschitz activation function and allowing the layer widths to grow to infinity at arbitrary relative rates. In the special case where all widths are proportional to a common scale parameter $n$ and there are L1L-1 hidden layers, we obtain convergence rates of order $n^{-({1}/{6})^{L-1}+\epsilon}$, for any $\epsilon>0$.
@@ 1 Introduction
Apollonio et al. (2025a) non-Lip FDD $W_{1}$ (w.r.t. 2-norm) & Convex distance $n^{-\tfrac{1}{2}}$
Trevisan (2023) Lip FDD $W_{p}$ (w.r.t. $p$-norm) for $p\in[1,\infty)$ $n^{-\tfrac{1}{2}}$ & $n^{-1}$ (full-rank limiting covariance)
@@
Celli and Peccati (2025) Poly bounded FDD TV & $W_{2}$ (w.r.t. 2-norm) $n^{-1}$ (full-rank limiting covariance)
@@
where $\sigma$ is applied coordinate-wise. Typically, each layer also includes Gaussian bias parameters; however, setting them to zero does not affect our results and is therefore omitted—see Remark 1.5. Given training data, the weights of the neural network are chosen by minimizing a loss function using a gradient descent algorithm with a random initialization, where the weights $W_{i,j}^{(\ell)}$ are chosen to be centered, identically distributed, and independent across $i,j,\ell$. In the wide regime where $n_{i}$ is large for all $i=1,\ldots,L-1$, it is known that the neural network at such an initialization is close to a Gaussian process. This was first observed by Neal (1996), who showed the convergence of a single hidden-layer neural network at initialization to a Gaussian. The Gaussian limit in the single layer case is fairly straightforward because $F^{(2)}$ can be written as a sum of independent functions, and so classical techniques for sums of independent random variables apply. Subsequent works such as Lee et al. (2018) and de G. Matthews et al. (2018) provided heuristic arguments and empirical evidence suggesting that deep neural networks with multiple layers similarly exhibit Gaussian behavior as their widths grow. More recently, Hanin (2023) rigorously established asymptotic convergence results for deep networks, further supporting the Gaussian behavior in the infinite-width regime. Our main result establishes Gaussian approximation bounds, in the Wasserstein-1 distance, between the finite-dimensional distributions (FDDs) of wide neural networks, and these Gaussian process limits, under general independent weight distributions satisfying mild moment conditions, and assuming a Lipschitz activation function.
@@
Apart from their intrinsic probabilistic interest, there are several compelling reasons to study deep neural networks (DNNs) with non-Gaussian weights. In practice, weights are typically initialized randomly—often with independent and identically-distributed entries—and subsequently optimized using algorithms like stochastic gradient descent (SGD) to minimize a loss function over a given dataset. While Gaussian initializations are widely used, other schemes such as uniform initialization (Glorot and Bengio, 2010; Golikov and Yang, 2022) and Bernoulli distributions, particularly in quantized networks (Li et al., 2017), are also common. Moreover, in transfer learning settings where pre-trained models are fine-tuned on new tasks, the initial weight distribution is often far from Gaussian. This motivates the study of DNN behavior under more general random weight initializations. Furthermore, Jacot et al. (2018) showed that in the so-called lazy training regime, infinitely wide DNNs exhibit deterministic training dynamics governed by the Neural Tangent Kernel (NTK), and their fluctuations converge to Gaussian processes—emphasizing how the choice of initialization fundamentally shapes both the limiting kernel and the resulting generalization behavior (Carvalho et al., 2023). Finally, several works have explored DNNs with weights drawn from heavy-tailed or stable distributions, which may exhibit infinite variance (Der and Lee, 2005; Favaro et al., 2023; Bordino et al., 2023; Soto, 2024; Jung et al., 2023; Loría and Bhadra, 2023). These studies are motivated by empirical advantages observed in such models and have established that, in the infinite-width limit, network outputs may converge to stable processes rather than Gaussian processes.
@@
A recent and very thorough account of this literature is given in Celli and Peccati (2025, Section 1.7), but we give a quick review to contextualize our results. A summary of quantitative bounds for deep neural networks is provided in Table 1. For shallow NNs, some functional results include Eldan et al. (2021), Klukowski (2022), and Cammarota et al. (2024), all with fairly restrictive assumptions on the activations or weight distributions, and typically in weak $L^{p}$-type metrics at the functional level. For deep NNs, rates of convergence for FDDs assuming Gaussian weights have been given in Apollonio et al. (2025a), Basteri and Trevisan (2024), Bordino et al. (2024), Celli and Peccati (2025), Favaro et al. (2025), and Trevisan (2023), under a variety of activations and metrics. At the functional level, assuming Gaussian weights, Favaro et al. (2025) derive rates of convergence with respect to Sobolev topologies in the Wasserstein-$2$ norm, and, assuming the activation is infinitely smooth, with respect to the supremum topology. Assuming general weights, process convergence with respect to the supremum topology was shown by Hanin (2023), and rates of convergence in the same topology in the Wasserstein-$1$ norm are given in Balasubramanian et al. (2024).
@@ 1.1 Main result
In this work, we provide a bound on the Wasserstein-$1$ norm between the FDDs of $F^{(L)}$ and their Gaussian limit assuming a general weight distribution and a Lipschitz activation function. The bound can be made explicit in all parameters, is independent of spectral properties of the limiting covariance, and tends to zero as the widths tend to infinity in any way, for fixed $L$. This is the first such bound appearing in the literature; see also Table 1.
@@
Given this, Hanin (2023) showed that for an absolutely continuous activation with polynomially-bounded derivative and centered independent and identically-distributed weights $W_{ij}^{(\ell)}$ with all absolute moments $p\geqslant 1$ bounded uniformly of order $n_{\ell}^{-p/2}$, the neural network at initialization converges to a Gaussian field which is defined recursively via intermediate fields $G^{(\ell)}$, $\ell=1,\ldots,L-1$ as follows. We first set $G^{(1)}=F^{(1)}$, and, given we have defined the distribution of $G^{(\ell)}$, we let $G^{(\ell+1)}$ be a centered Gaussian random field with covariance
$C_{ij}^{(\ell+1)}(x,y)=\delta_{ij}\biggl(c_{w}^{(\ell)}\mathbbm{E}\Bigl[\sigma\bigl(G^{(\ell)}_{1}(x)\bigr)\sigma\bigl(G^{(\ell)}_{1}(y)\bigr)\Bigr]\biggr),$ (1.3)
where $\delta_{ij}$ denotes the Kronecker delta, and $G_{1}^{(\ell)}$ is the first component of $G^{(\ell)}$. Note that the covariance function of $F^{(L)}$, calculated from (1.1) and (1.2), follows the same recursion, but with $G$ replaced by $F$, so that the coordinates of $F^{(\ell)}$ are uncorrelated, though not independent. An important component of most proofs in the deep regime is that if $F^{(\ell)}$ is approximately Gaussian, then the coordinates should be “close” to independent.
@@
$\mathsf{d}_{1}\bigl(F^{(L)}(\chi),G^{(L)}(\chi)\bigr):=\sup_{\zeta\in\mathcal{W}_{n_{L}\times s}}\bigl\lvert\mathbbm{E}\bigl[\zeta(F^{(L)}(\chi))\bigr]-\mathbbm{E}\bigl[\zeta(G^{(L)}(\chi))\bigr]\bigr\rvert$,
@@
Let $F^{(L)}$ be the DNN defined at (1.1) with centered weights $W_{i,j}^{(\ell)}$ satisfying (1.2) which are independent across $i,j,\ell$ with identically distributed rows: $(W_{i,k}^{(\ell)})_{k=1}^{n_{\ell}}\stackrel{{\scriptstyle d}}{{=}}(W_{j,k}^{(\ell)})_{k=1}^{n_{\ell}}$, and with a Lipschitz activation $\sigma$. Assume further that there is $p>2$ and constants $c_{2p}^{(\ell)}\geqslant 1$ for $\ell=0,\ldots,L-2$ such that $\mathbbm{E}\bigl[(W_{ij}^{(\ell)})^{2p}\bigr]\leqslant c_{2p}^{(\ell)}/n_{\ell}^{p}$ and a constant $c_{3}^{(L-1)}$ such that $\mathbbm{E}\bigl[\lvert W_{ij}^{(L-1)}\rvert^{3}\bigr]\leqslant c_{3}^{(L-1)}n_{L-1}^{-3/2}$. Then for $\chi=(x_{1},\ldots,x_{s})\in\mathcal{M}^{s}$, there is a constant $\mathtt{C}$ depending on $\sigma$, $p$, $L$, $\chi$, $c_{2p}^{(\ell)}$, $\ell=0,\ldots,L-2$ and $c_{3}^{(L-1)}$ such that
where $G^{(L)}$ is the Gaussian process defined by the covariance recursion (1.3).
@@
$n_{L}^{1/3}\sum_{m=1}^{L-1}\frac{1}{n_{m}^{\left(1/6\right)^{L-m}-\delta}},$
@@
$n_{L}^{\varphi}\sum_{m=1}^{L-1}n_{m}^{-\frac{1}{2}},$ (1.4)
for some $\varphi>0$. Our worse power stems from initially working in an integral probability metric $\mathsf{d}_{3}$ defined at (2.1), which is weaker than $\mathsf{d}_{1}$, but can be easily used with Stein’s method and does not require any assumptions on the covariance matrix of the limiting Gaussian vector $G^{(L)}(\chi)$ (see the forthcoming Remark 1.4), without which the rate may in truth be worse than the classical setting. We then use a smoothing argument to get back to $\mathsf{d}_{1}$ which introduces a factor of $1/3$ to the power. (There is a further factor of $1/2$ introduced for reasons too technical to explain in a remark, but it stems from (2.14).) The argument is inductive in the layers, and the smoothing is performed at each step, which leads to the power of (Lm)(L-m) in the factor of $1/6$ in the power. Similar to Balasubramanian et al. (2024, Section 6.2), the rate could be somewhat improved by assuming the activation function has three bounded derivatives. The form of the rate would still have the power of (Lm)(L-m) in the exponent, but now with a base of $(1/2)$ (stemming from (2.14)) rather than $(1/6)$.
@@
It is worth noting that our result requires no assumption on the limiting covariance of $G^{(L-1)}(\chi)$. As previously mentioned, this is because we first work in the weaker $\mathsf{d}_{3}$ metric; see Theorem 2.10. The potential degeneracy of the covariance figures prominently in Gaussian approximation results for DNNs, and many rates depend on the eigenvalues of the limiting covariance, including demanding they are all strictly positive (see Table 1). This is because when the weights are Gaussian, $F^{(L)}$ is conditionally Gaussian given $F^{(L-1)}$, and comparing a conditional Gaussian to $G^{(L)}$ requires comparing the (conditional) covariances; see Basteri and Trevisan (2024, (2.7),(2.1), and Lemma 3.4) and Favaro et al. (2025, Propositions 5.11 and 5.12); which involves the spectrum of covariance matrices. Applying such results may require additional hypotheses on the NN, and $\sigma$ especially. In contrast, our constants do not involve eigenvalues of the covariance. Revisiting the discussion of the best achievable rates of Remark 1.3, since DNNs do not fit neatly into the setting of the classical central limit theorem, the rate (1.4) may not be achievable without additional assumptions.
@@
As is typical with approximation results for DNNs in transportation metrics, the Gaussian biases that usually appear in the definition (1.1) of the NN do not affect the error in our Theorem 1.1. See Balasubramanian et al. (2024, Proof of Theorem 1.2) for the argument in the more challenging infinite-dimensional setting. As discussed in Hanin (2023, Section 1.4), the biases must be Gaussian to have a Gaussian limit.
Remark 1.6.
Recent work by Lee et al. (2023) and Apollonio et al. (2025b) shows that the standard assumption of independent weights can be relaxed by introducing per-node variance mixtures, allowing the outgoing weights from a neuron to share dependence through a latent variance parameter while still admitting a meaningful infinite-width limit. In particular, when the latent mixing distribution is non-trivial, the resulting limit is no longer a single Gaussian process but a mixture of Gaussian processes, exhibiting potentially heavy-tailed behavior. In our setting, one could in principle adopt a similar mixture-based relaxation, though this would substantially complicate the analytical tractability and alter the nature of the limiting kernel object, and consequently, the associated convergence rates. A detailed investigation is left as future work.
@@ 2 The proof
$\mathsf{d}_{3}\bigl(F^{(L)}(\chi),G^{(L)}(\chi)\bigr):=\sup_{\zeta\in\mathcal{F}_{n_{L}\times s}}\bigl\lvert\mathbbm{E}\bigl[\zeta(F^{(L)}(\chi))\bigr]-\mathbbm{E}\bigl[\zeta(G^{(L)}(\chi))\bigr]\bigr\rvert$, (2.1)
@@
where $\partial_{(\cdot)}$ denotes the partial derivative with respect to the coordinate(s) given by $(\cdot)$. We then move back to the Wasserstein distance using standard smoothing arguments; see Lemma 2.11.
For $F^{(\ell)}(\chi)=(F_{1}^{(\ell)}(\chi),\ldots,F_{n_{\ell}}^{(\ell)}(\chi))\in\mathbbm{R}^{n_{\ell}\times s}$ and any $1\leqslant d\leqslant n_{\ell}$, we write
$F_{[d]}^{(\ell)}(\chi):=(F^{(\ell)}_{1}(\chi),\ldots,F^{(\ell)}_{d}(\chi))\in\mathbbm{R}^{d\times s}.$
Our first step is to look at the distance between an intermediate field. Let $(\widetilde{W}^{(\ell)})_{\ell=0}^{L-1}$ be independent Gaussian weight matrices having the same variances as $(W^{(\ell)})_{\ell=0}^{L-1}$ given at (1.2), with the two sequences of weight matrices independent so that, e.g., $\widetilde{W}^{(\ell-1)}$ is independent of $F^{(\ell-1)}$. Let $\widetilde{F}^{(\ell)}=\widetilde{W}^{(\ell-1)}\sigma(F^{(\ell-1)})$. Then the triangle inequality implies for an $d=1,\ldots,n_{L}$, we have